# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

### v3

@@ Abstract.
A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first efficient, physically based mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with hydrometeor electrification processes. Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and efficient approach for graphical applications. By incorporating key physical principles, it effectively links cloud formation, electrification, and lightning generation. The simulation also encompasses various thunderstorm types and their corresponding lightning activities. For more details, see the dynamic video: Thunderscapes Video.
@@
ccs: Computing methodologies Physical simulation
@@ 1. INTRODUCTION
Thunderstorms represent the wild power of nature and are a common atmospheric element in visual effects (VFX) industry. Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure 2), demonstrate the importance of realistic atmospheric effects. Modern applications demand tools for efficient and realistic simulation of thunderstorms. However, current general purpose VFX software, such as Houdini and Maya, lacks standardized toolkits specifically designed for creating atmospheric phenomena like thunderstorms.
(a) Thunderstorm from *Ghost of Tsushima*. ©2020 Sony Interactive Entertainment LLC.
[…]
A scenic image from the video game Ghost of Tsushima showing dramatic landscapes and warrior scenes.
(b) Thunderstorm from *Horizon Forbidden West: Burning Shores*. ©2023 Sony Interactive Entertainment Europe.
[…]
A scene from the video game Horizon Forbidden West: Burning Shores, showcasing futuristic landscapes and action.
Figure 2. Comparison of atmospheric visual styles between *Ghost of Tsushima* and *Horizon Forbidden West: Burning Shores*.
Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes (Hädrich et al., 2020; Herrera et al., 2021; Amador Herrera et al., 2024). However, the critical role of atmospheric electrification in cloud development remains underexplored. This study aims to address this gap by integrating the electrification process, thus enhancing the consistency and realism of phenomena observed during thunderstorm events.
This paper presents an efficient, physically based method for simulating thunderstorm development. By coupling cloud microphysics with atmospheric hydrometeor electrification, our approach captures essential processes that contribute to the consistent formation of mesoscale thunderstorm phenomena.
The key contributions include:
@@
    We present a comprehensive framework for computing thunderstorm microphysics in the atmosphere, incorporating the transport of cloud hydrometeors and their electrification processes.
@@
    We introduce a lightweight parameterization that enables users to efficiently simulate various types of thunderstorms, along with their corresponding lightning effects.
@@
    We validate our simulation using different meteorological datasets to ensure the generation of physically credible and visually accurate atmospheric effects.
@@ 2. RELATED WORK
2.1. Simulating Thunderstorms in Computer Graphics
The simulation of atmospheric phenomena such as thunderstorms has been extensively explored through a variety of computational methods. Webanck et al. (Webanck et al., 2018) proposed a procedural approach for generating cloudscapes, while Miyazaki et al. (Miyazaki et al., 2002) simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations. Ferreira et al. (Ferreira Barbosa et al., 2015) and Zhang et al. (Zhang et al., 2020) utilized position-based fluids (PBF) for adaptive cloud simulations. Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret (Goswami and Neyret, 2017), focus on real-time simulations of convective clouds. Additionally, Vimont et al. (Vimont et al., 2020) proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes.
Some studies have specifically focused on the simulation of volcanic cloud dynamics. Lastic et al. (Lastic et al., 2022) employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al. (Pretorius et al., 2024) integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes.
@@ 2.1. Simulating Thunderstorms in Computer Graphics
Recently, more sophisticated microphysical schemes from atmospheric science have been incorporated into computer graphics research. Garcia-Dorado et al. (Garcia-Dorado et al., 2017) and Hädrich et al. (Hädrich et al., 2020) adopted the classic Kessler warm cloud microphysics scheme (Kessler, 1969) in cloud simulations. Herrera et al. (Herrera et al., 2021) extended cloud simulations to include multiphase cloud dynamics. Amador Herrera et al. (Amador Herrera et al., 2024) developed a framework to simulate hurricane and tornado dynamics.
@@ 2.2. Thunderstorm models in atmospheric sciences
Thunderstorm microphysics and electrification have been widely discussed topics in the field of atmospheric science.
Kessler (Kessler, 1969) proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
@@
Our work focuses on the efficient simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS). According to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, these types include single cell, multi-cell, squall line, and supercell thunderstorms. Additionally, our simulation includes common phenomena generated by thundercloud electrification, such as cloud-to-ground (CG) and intra-cloud (IC) lightning [^2] https://www.nssl.noaa.gov/education/svrwx101/lightning/types/.
Figure 3. Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process. The details will be further discussed in Chapter 4. (I) Blue frame: the transport of water substances during thundercloud development; (II)Orange frame: the accumulation of charge density due to the collision-coalescence of hydrometeors; (III) Red frame: the static charge, once it exceeds a given electric field threshold, triggers the lightning discharge process; (IV) Blue arrow: the neutralization of hydrometeors following the lightning flash.
@@
3. OVERVIEW
The primary motivation of our approach is to create consistent atmospheric phenomena during thunderstorms, particularly focusing on thundercloud development and dissipation, along with lightning flashes resulting from the thundercloud electrification processes,as illustrated in Figure 3. By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, efficient weather simulations suitable for graphical applications.
Our model integrates key physical principles to simulate thunderstorm behavior at a mesoscale level. The thunderstorm microphysics component models atmospheric dynamics by coupling a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processing. This approach ensures that the microphysical processes driving cloud formation and growth are consistently linked to the electrification necessary for lightning generation.
The key atmospheric quantities driving our model include vapor, cloud water, ice, precipitated rain , precipitated snow , static charge (electric charge distribution within the thundercloud, crucial for lightning formation), and lightning (the flash itself, including both its onset and development).
The validation of our model involves presenting spatially simulated thundercloud structures alongside meteorological characteristics and comparing the temporally simulated results with real-world weather data. This includes analyzing cloud fraction profiles to assess cloud formation and structure, tracking the evolution of cloud coverage against real-time data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and evaluating the temporal variation in lightning flash rates.
Figure 4. Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
[…]
A schematic representation of the thunderstorm microphysics scheme, coupling the extended warm cloud microphysics with the electrification process of hydrometeors.
@@ 4. METHODOLOGY
Our microphysics model, shown in Figure 4, illustrates the interactions among hydrometeor phase transitions and the electrification mechanisms that lead to lightning. Key processes include the condensation of water vapor into droplets and ice crystals, which subsequently form precipitation through autoconversion and accretion. Evaporation recycles hydrometeors back into the vapor phase, while collisions and coalescence facilitate charge separation. Lightning occurs when the electric field strength exceeds a critical threshold, resulting in the discharge and redistribution of accumulated charge. This feedback loop provides a conceptual framework for understanding thunderstorm development within a mesoscale convective system.
@@ 4.1. Cloud Microphysics
The fundamental warm-cloud microphysics equations describe the evolution of potential temperature, water vapor, cloud condensate, and precipitation:
(1) $\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$
(2) $\frac{Dq_{v}}{Dt}=-COND+DIFF$
(3) $\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$
(4) $\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$
@@
Here, $\theta$ represents potential temperature, $L_{v}$ is the latent heat of condensation or evaporation, and $c_{p}$ is the specific heat capacity at constant pressure. The variables $q_{v}$, $q_{c}$, and $q_{p}$ denote the mixing ratios of water vapor, cloud condensate, and precipitation, respectively. The terms $COND$, $DIFF$, $AUTC$, and $ACCR$ represent the condensation rate, diffusional growth rate of precipitation, autoconversion rate, and accretion rate, respectively.
@@
We adopt the equilibrium approach proposed by Grabowski, incorporating a temperature-dependent factor $\alpha$ to differentiate warm and cold cloud microphysics. Warm clouds dominate above $0^{\circ}\mathrm{C}$ and cold clouds below $-20^{\circ}\mathrm{C}$, with a linear interpolation for intermediate temperatures:
(5) $q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$
(6) $q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$
(7) $q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$
Here, $q_{vs}$ represents the total vapor saturation, combining vapor saturation over water ($q_{vw}$) and ice ($q_{vi}$); $q_{w}$ and $q_{i}$ denote cloud water and ice, expressed as fractions of the cloud-ice mixture ($q_{c}$); $q_{r}$ and $q_{s}$ represent rain and snow, derived as fractions of the rain-snow mixture ($q_{p}$).
The transport equations for rain and snow processes include separate contributions from diffusional growth ($DIFF$), autoconversion ($AUTC$), and accretion ($ACCR$) rates:
(8) $DIFF=DIFF_{r}+DIFF_{s}$
(9) $AUTC=AUTC_{r}+AUTC_{s}$
(10) $ACCR=ACCR_{r}+ACCR_{s}$
The terminal velocity $v_{p}$ of precipitation is modeled as a weighted combination of rain ($v_{r}$) and snow ($v_{s}$) velocities, with typical values $v_{r}=-10\,\mathrm{m/s}$ for rain and $v_{s}=-2\,\mathrm{m/s}$:
(11) $v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$
@@
The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$(Yau and Rogers, 1996):
(12) $q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$
@@
Here, $p$ represents the pressure, and $T$ denotes the temperature. The saturation vapor is modeled as an exponential distribution for both liquid water and ice.
The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions (Dudhia, 1989),where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
(13) $DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$
@@
The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation (Kessler, 1995; Lin et al., 1983):
(14) $AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$
@@
The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles (Morrison et al., 2015):
(15) $ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$
4.2. Electrification

### v4

@@ Abstract.
A Mesoscale Convective System (MCS) is a collection of thunderstorms operating as a unified system, showcasing nature’s untamed power. They represent a phenomenon widely referenced in both the natural sciences and the visual effects (VFX) industries.However, in computer graphics, visually accurate simulation of MCS dynamics remains a significant challenge due to the inherent complexity of atmospheric microphysical processes.To achieve a high level of visual quality while ensuring practical performance, we introduce Thunderscapes, the first physically based simulation framework for visually realistic MCS tailored to graphical applications.Our model integrates mesoscale cloud microphysics with hydrometeor electrification processes to simulate thunderstorm development and lightning flashes. By capturing various thunderstorm types and their associated lightning activities, Thunderscapes demonstrates the versatility and physical accuracy of the proposed approach.
@@
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3728312journal: PACMCGITjournalvolume: 8journalnumber: 1article: 7publicationmonth: 5ccs: Computing methodologies Physical simulation
@@ 1. INTRODUCTION
Thunderstorms represent the wild power of nature and are a common atmospheric element in the visual effects (VFX) industry. Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima, demonstrate the importance of realistic atmospheric effects. Modern applications demand tools for efficient and realistic simulation of thunderstorms.
However, current general-purpose VFX software, such as Houdini and Maya, relies on traditional Eulerian fluid dynamics (e.g., stable fluids (Stam, 1999)) to simulate phenomena like smoke or explosions. They fall short in capturing the inherent complexity of atmospheric microphysical processes, which are essential for accurately simulating the dynamics of complex thunderstorm systems, such as MCS, and for producing consistent, high-quality results.
This paper introduces Thunderscapes, a physically based simulation framework for visually realistic MCS. By incorporating a Grabowski-style cloud microphysics model (Grabowski, 1998) alongside hydrometeor electrification processes, our approach delivers high visual fidelity while maintaining practical performance levels, capturing the core dynamics of thunderstorm evolution and lightning activities within an MCS. Furthermore, the framework integrates seamlessly with Houdini, a widely used tool in the VFX industry, through intuitive and lightweight parameters, allowing artists to utilize its advanced simulation features without requiring expertise in meteorology or knowledge of supercomputing.
The key contributions of this work are as follows:
@@
    We develop a realistic, physically based simulation framework for MCS, integrating mesoscale cloud microphysics with hydrometeor electrification processes to simulate diverse thunderstorm phenomena.
@@
    We present a lightweight set of parameters that allows artists to intuitively create thunderstorm animations within general-purpose VFX software.
@@
    We validate the framework using diverse meteorological datasets, demonstrating its capability to generate visually realistic and physically consistent atmospheric effects.
@@ 2. RELATED WORK
2.1. Simulating Thunderstorm in Computer Graphics
The visual simulation of atmospheric phenomena such as convective clouds has been extensively explored through a variety of computational methods. Webanck et al. (Webanck et al., 2018) proposed a procedural approach for generating cloudscapes, while Miyazaki et al. (Miyazaki et al., 2002) simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations. Ferreira et al. (Ferreira Barbosa et al., 2015) and Zhang et al. (Zhang et al., 2020) utilized position-based fluids (PBF) for adaptive cloud simulations. Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret (Goswami and Neyret, 2017), focus on real-time simulations of convective clouds. Additionally, Vimont et al. (Vimont et al., 2020) proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes. However, these approaches lack a detailed focus on cloud microphysics, limiting their ability to produce realistic and complex cloudscapes.
@@ 2.1. Simulating Thunderstorm in Computer Graphics
Recently, more sophisticated microphysical schemes from atmospheric science have been incorporated into computer graphics research. Garcia-Dorado et al. (Garcia-Dorado et al., 2017) and Hädrich et al. (Hädrich et al., 2020) adopted the classic Kessler warm cloud microphysics scheme (Kessler, 1969) in cloud simulations. Herrera et al. (Herrera et al., 2021) extended cloud simulations to include multiphase cloud dynamics. Amador Herrera et al. (Amador Herrera et al., 2024) developed a framework to simulate hurricane and tornado dynamics.However, these methods overlook the hydrometeor electrification process, which limits their ability to simulate consistent and realistic thunderstorms with associated lightning phenomena.
@@ 2.2. Thunderstorm models in atmospheric sciences
Thunderstorm microphysics and electrification have been widely discussed topics in the field of atmospheric science.Kessler (Kessler, 1969) proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
@@
3. OVERVIEW
The primary motivation for our approach is to visually capture the realistic development of MCS. We propose a physically based simulation framework that is intuitive for artists, enabling the creation of visually realistic thunderstorm animations while managing physical complexity and ensuring practical performance.
Simulating MCS is challenging due to the complex microphysical and electrostatic dynamics and the interactions among hydrometeors (ice, snow, and rain) that induce charge imbalances. Traditional meteorological models capture these phenomena with high accuracy but rely on numerous variables and detailed phase transitions(Barthe and Pinty, 2007), making them computationally expensive and difficult for non-experts to manage.For example,the conventional warm-cloud model typically involves three or four phase transitions and works well for ordinary cumulus clouds, whereas modeling complex thunderstorms generally requires a combined warm-cold cloud scheme with more than ten phase transitions (e.g., melting and riming (Rutledge and Hobbs, 1984)). Although such detailed modeling improves numerical weather prediction, its prohibitively long computation times necessitate the use of supercomputers.
In contrast,our Grabowski-style framework extends the basic warm-cloud model with a temperature interpolation scheme to incorporate cold-cloud processes, capturing the essential material transformations during thunderstorm development with fewer phase transitions. This simplification sacrifices some microscopic detail yet preserves a physically accurate macroscopic evolution, reduces the number of user-controlled variables, and ensures practical performance.
Building on these advantages, we develop a comprehensive framework to describe the essential microphysical processes of hydrometeor phase transitions and electrification.Our model employs six key parameters to characterize the atmospheric state: the vapor field $q_{v}$, cloud field $q_{c}$, precipitation field $q_{p}$, temperature field $\theta$, charge density field $\rho$, and velocity field $\mathbf{u}$.Our simulation framework, as illustrated in Figure 2, consists of the following components:(I) Cloud microphysics, describing hydrometeor phase transitions such as condensation, evaporation, and precipitation; (II) Hydrometeor electrification, modeling charge accumulation during hydrometeor interactions; (III) Lightning discharge, computing charge neutralization as lightning channels propagate; (IV) Atmospheric background, accounting for buoyancy forces arising from temperature gradients; (V) Fluid dynamics, driving field evolution with an Eulerian solver for MCS development.
Figure 2. Schematic illustration of our MCS life cycle, encompassing the interplay and feedback among key modules. Detailed explanations are provided in Section 4.(I) Cloud Microphysics: Describes the phase transitions of hydrometeors during MCS development.(II) Hydrometeor Electrification: Explains the accumulation of charge density resulting from the collision and coalescence of hydrometeors.(III) Lightning Discharge: Highlights the process of lightning channel formation and propagation, triggered when static charge exceeds a predefined electric field threshold, leading to the neutralization of hydrometeors along the lightning channel.(IV) Atmospheric Background: Introduces vertical buoyancy forces to the system driven by temperature variations caused by hydrometeor phase transitions.(V) Fluid Dynamics: Drives the system’s temporal evolution and introduces forces influencing the dynamics of the MCS life cycle.
@@ 4. METHODOLOGY
Our microphysics model, depicted in Figure 2, illustrates the interactions among hydrometeor phase transitions and electrification mechanisms. Key processes include the condensation of water vapor into droplets and ice crystals, which subsequently form precipitation through autoconversion and accretion. Evaporation recycles hydrometeors back into the vapor phase, as detailed in Section 4.1. Collisions and coalescence facilitate charge separation, leading to lightning when the electric field strength exceeds a critical threshold, as described in Section 4.2. The buoyancy forces generated by atmospheric temperature variations are discussed in Section 4.3, while the system’s dynamics updates are outlined in Section 4.4. This feedback loop provides a comprehensive framework for understanding thunderstorm development within an MCS.
@@ 4.1. Cloud Microphysics
The fundamental warm-cloud microphysics equations govern the evolution of potential temperature, water vapor, cloud condensate, and precipitation. These equations are formulated using the material derivative (Kundu et al., 2012):
(1) $\frac{\mathrm{D}\phi}{\mathrm{D}t}=\frac{\partial\phi}{\partial t}+\boldsymbol{u}\cdot\nabla\phi,$
which provides the foundation for expressing the microphysics equations:
(2) $\frac{\mathrm{D}\theta}{\mathrm{D}t}=\frac{L_{v}\theta}{C_{p}T}(C_{C}-D_{P}),$
(3) $\frac{\mathrm{D}q_{v}}{\mathrm{D}t}=-C_{C}+D_{P},$
(4) $\frac{\mathrm{D}q_{c}}{\mathrm{D}t}=C_{C}-A_{C}-G_{C},$
(5) $\frac{\mathrm{D}q_{p}}{\mathrm{D}t}=A_{C}+G_{C}-D_{P},$
@@
where $\theta$ represents the potential temperature, $L_{V}$ is the latent heat of condensation or evaporation, $T$ is the absolute temperature, and $C_{p}$ is the specific heat capacity at constant pressure. Additionally, $q_{v}$, $q_{c}$, and $q_{p}$ denote the mixing ratios of water vapor, cloud condensate, and precipitation, respectively. The terms $C_{C}$, $D_{P}$, $A_{C}$, and $G_{C}$ correspond to the condensation rate, diffusional growth rate of precipitation, autoconversion rate, and accretion rate.
@@
We adopt the equilibrium approach proposed by Grabowski (Grabowski, 1998), incorporating a temperature-dependent factor $\alpha$ to differentiate warm and cold cloud microphysics. The parameter $\alpha$ varies linearly with temperature: warm clouds dominate above $0^{\circ}\mathrm{C}$, cold clouds dominate below $-20^{\circ}\mathrm{C}$. The unified relationship for vapor saturation, cloud, and precipitation components is expressed as:
(6) $q_{mc}=\alpha q_{wc}+(1-\alpha)q_{cc},$
where $q_{mc}$ represents the total quantity of a variable $q_{vs},q_{c},q_{p}$, combining its warm-phase component $q_{wc}$ and cold-phase component $q_{cc}$. Specifically, for vapor saturation, $q_{wc}=q_{vw}$ (saturation over water) and $q_{cc}=q_{vi}$ (saturation over ice); for cloud condensate, $q_{wc}=q_{w}$ (cloud water) and $q_{cc}=q_{i}$ (cloud ice); and for precipitation, $q_{wc}=q_{r}$ (rain) and $q_{cc}=q_{s}$ (snow).
@@
Building on this framework, we parameterize diverse microphysical processes, unified by the temperature-dependent factor $\alpha$. The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice(Yau and Rogers, 1996):
(7) $q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right],$
@@
where $p$ represents the pressure, and $T$ denotes the temperature. The saturation vapor is modeled as an exponential distribution for both liquid water and ice. The condensation rate $C_{C}$ governs cloud formation from water vapor, integrating water and ice contributions (Dudhia, 1989), where $\beta_{w},\beta_{i}$ are phase-specific coefficients:
(8) $C_{C}=\beta_{w}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{i}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}.$
@@
The autoconversion rate $A_{C}$ describes cloud condensate aggregation into precipitation (Kessler, 1995; Lin et al., 1983):
(9) $A_{C}=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}.$
@@
The accretion rate $G_{C}$ models the collection of cloud condensate by precipitation particles (Morrison et al., 2015):
(10) $G_{C}=q_{c}\cdot q_{p}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right).$
The diffusional growth rate $D_{P}$ describes the conversion between precipitation and vapor through deposition and evaporation processes (Klemp and Wilhelmson, 1978), and is modeled as follows:
(11) $D_{P}=\alpha(1-\alpha)\frac{q_{vs}-q_{v}}{\hat{\rho}q_{vs}}\frac{\left(\hat{\rho}q_{p}\right)^{0.525}}{5.4\cdot 10^{5}+2.55\cdot 10^{8}\left(pq_{vs}\right)^{-1}},$
where $\hat{\rho}$ is the density of humid air, calculated in the same way as in (Hädrich et al., 2020), and $p$ represents the atmospheric pressure, which will be described in Section 4.3. The terminal velocity $\boldsymbol{u}_{p}$ of precipitation during deposition is modeled as a weighted combination of rain $\boldsymbol{u}_{r}$ and snow $\boldsymbol{u}_{s}$ velocities.Since precipitation particles quickly reach a uniform terminal velocity, we use typical values of $\boldsymbol{u}_{r}=-10\,\mathrm{m/s}$ for rain and $\boldsymbol{u}_{s}=-2\,\mathrm{m/s}$ for snow:
(12) $\boldsymbol{u}_{p}=\alpha\boldsymbol{u}_{r}+(1-\alpha)\boldsymbol{u}_{s},$
4.2. Hydrometeor Electrification